• Main INDEX
  • Monthly INDEX
  • PREV
  • NEXT
    Make New Entry, Make Followup Entry

    User name Saha

    Log entry time 16:03:42 on November10,2009

    Entry number 298760

    keyword=Shift Summary

    Shift Summary
    0800:  RHRS Magnets down,  liquid level low. Gary looking into problem
           MCC Injector work started
    0941:  Target heater control beam compensation is fixed by Jian-Ping   
    1030: Call MCC for Controlled access to fix magnets problems
    1600: Hall still in controlled access. Javier and Jack still in hall.